# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

### v3

@@ III-C Baselines and Results
III-D Summary
The LaMPilot-Bench dataset and the CARLA Leaderboard 1 benchmark provide a comprehensive evaluation framework for assessing the instruction-following capabilities and performance of LLM-based agents in AD. The experiments highlight the potential of LLMs in enhancing the decision-making capabilities of autonomous vehicles when combined with human feedback and domain-specific APIs. As LLMs continue to evolve and integrate with more modalities, we envision their significant contribution to the development of safe, efficient, and user-friendly autonomous vehicles. Future research should focus on addressing the challenges of real-time decision-making, reducing collision rates, and incorporating multi-modal inputs to further enhance the performance and robustness of LLM-based autonomous driving systems.
@@ III-D Summary
IV Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making
@@
IV-A Talk2Drive Framework
@@
IV-A1 Problem Statement
In Talk2Drive, we follow the concept we proposed in Sec. II. The model aims to translate verbal commands into executable control sequences for the vehicle. The cloud-based LLM acts as a translating function $f:\boldsymbol{I}\rightarrow\boldsymbol{P}$ that generates corresponding LMPs as the policy ($\boldsymbol{P}$) for maneuvers.
For personalization, we need human feedback on execution $F(I,P)$ to evaluate if the generated policy addresses the preferences of the driver or the passengers. The extra memory module (see IV-A5) includes instructions, LMPs, and their evaluation will allow to guide the LLM to learn preferences. Therefore, there are two stages in the Talk2Drive workflow:
@@ IV-A1 Problem Statement
[…]
IV-A2 Command Translation and Contextual Data Integration
$I\to\text{Could you drive more conservatively?}$? (10)
Simultaneously, LLMs access additional cloud-based real-time environment data, including weather updates, traffic conditions, and local traffic rules. In our product, LLMs can be informed by the weather information through Openweather API [openweather], the map information (such as road type and speed limits) through OpenStreetMap API [OpenStreetMap], and traffic information through TomTom API [tomtom]. We use a predefined structured language generator for both $C$ and $S$, and our system will supply the appropriate values (the red contents) to the generator based on the context information, as shown below:
$C\to$ (11)
$\left\{\begin{array}[]{l}\text{A vehicle in front of you is running at {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}38.0} km/h.}\\ \text{Your current speed is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}40.0} km/h.}\\ \text{The speed limit is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}60.0} km/h.}\\ \text{The weather is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}sunny}.}\\ \vdots\end{array}\right.$ (17)
IV-A3 Processing and Reasoning with LLMs
@@
$\left\{\begin{array}[]{l}\text{You are an autonomous vehicle withAdaptive cruise}\\ \text{control (ACC) and Lane Keeping Assist (LKA) always}\\ \text{enabled.}\\ \text{You are using Pure Pursuit Controller to do the waypoint}\\ \text{following.}\\ \vdots\\ \text{Here are some examples of how you need to react.}\\ \text{Query: You drive too aggressively.}\\ \text{Thought: The drivers think I drive too fast which looks}\\ \text{aggressive and the drivers do not ask me to change lanes,}\\ \text{so I need to slow down my speed.}\\ \text{Action: ...}\\ \vdots\end{array}\right.$ (32)
Additionally, the LLMs also access the memory module, a repository of historical interactions ($H$), to consider the human’s past behaviors and preferences. More details about the memory module are in Sec. IV-A5.
IV-A4 Actionable Code Generation
@@
$\left\{\begin{array}[]{l}\text{The driver wants me to drive more conservatively but}\\ \text{does not have any specific lane preference. Therefore, I}\\ \text{should reduce my speed and maintain my current lane.}\\ \end{array}\right.$ (38)
The generated LMPs ($P$) are then sent back from the cloud to the vehicle’s ECU, where they are executed. Additionally, we also set two kinds of safety checks for the generated LMPs. The first check involves verifying the format of the generated code. If the code does not conform to a valid format, our system will not execute or respond to the code. The second safety check is parameter verification, which assesses the feasibility and safety of the parameters within the context of the given policy, preventing the execution of potentially risky policies. For example, if the generated code instructs our agent to move to the left lane while the ego agent is already in the leftmost lane, our system will identify this abnormality and prevent the execution.
IV-A5 Memory Module and Personalization
As described in Sec. II, the memory module is used to store the historical interactions ($H$) between humans and vehicles to provide a personalized driving experience. One important difference to note about this specific system, compared to the general framework, is that within this system, human evaluations $F$ are archived within the historical memory, and the history memory $H$ includes three key linked components: the past commands ($I$) issued by the human, the corresponding LMP ($P$) generated by the LLMs in response, and the humans’ subjective evaluations or feedback ($F$) on the system’s performance for each command. Each interaction between the human and the vehicle is recorded and saved into a memory module in a text format within the ECU. This historical data in the memory module is updated after every trip. The example contents in the memory module are shown in Eq. 39:
@@ IV-A5 Memory Module and Personalization
$\left\{\begin{array}[]{l}\text{Apart from the requirements I provided before. Here I}\\ \text{will provide you with the history dialogues between the}\\ \text{driver and the vehicle. You will need to learn what}\\ \text{drivers' wants and needs are.}\\ \text{For example, if the history driver's command is ``You}\\ \text{drive too conservatively.''}\\ \text{The history output action is : ...}\\ \text{After the trip, the driver's feedback is ``A little bit too}\\ \text{fast for me.''}\\ \text{Then the next time your action should adjust according}\\ \text{to the driver's feedback, which is ...}\\ \vdots\\ \text{The history command, action, and driver's feedback are:}\\ \text{Command: I'm on my way to the urgent care.}\\ \text{Action: ...}\\ \text{Evaluation: A little bit too fast.}\\ \vdots\end{array}\right.$ (57)
Figure 5: The flowchart of Talk2Drive. After the speech recognition module detects the keyword “command,” the inputs ($I,C,S,H$) are sent to the LLM. Then, the LLM generates corresponding LMPs to be executed by the ECU. If the speech recognition module detects the keyword “evaluate,” the system receives human feedback ($F$), and both $F$ and its corresponding $I$ and $P$ are updated in the memory module.
@@
IV-B Autonomous Driving System
IV-C Experiment Setup
IV-C1 Scenarios Setup and Participants
Our experimental trials include three different scenarios: highway, intersection, and parking lot.  [^2] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. The field experiments for the highway and intersection scenarios are conducted at a proving ground in Columbus, IN, USA, to validate the efficacy of the proposed Talk2Drive framework. The highway scenario involves a three-lane highway, while the intersection contains a two-way junction. Additionally, the parking lot scenario is evaluated at the North Stadium Parking Lot in West Lafayette, IN, USA. The overview visualization and statistics of the test tracks are shown in Fig. 6, while the visualization for the experiments is shown in Fig. 8. In our study, we recruit a diverse range of drivers, including individuals of different genders (61.4% male, 28.6% female), ages (mean=26.71, std=4.11), and driving experiences(mean=6.79, std=5.08). All participants have valid driving licenses.
IV-C2 Input Instructions
@@
Given that conventionally indirect strategies primarily modify the politeness level of command of direct strategies, we instead divide the non-conventionally indirect requests into two categories based on the strength of the hints. We define our levels of directness in Tab. V as our way to classify varying degrees of implicitness in the commands. To gather data, we request our test drivers to generate commands based on their normal speaking preferences and subsequently categorize their commands into our defined levels. Examples of the commands generated are also presented in Tab. V.
TABLE V: Level of Command Directness and Examples
@@
TABLE VI: Driving Performance Validation
@@
IV-C3 Evaluation Metrics
Our evaluation framework for autonomous vehicles includes driving performance, time efficiency, and personalization. We analyze Talk2Drive’s driving performance in terms of safety, using Time to Collision ($\tau$) and speed variance ($\sigma^{2}$), and comfort, using mean absolute acceleration ($|\bar{A}|$) and mean absolute jerk ($|\bar{J}|$). Time efficiency is measured by the LLM latency ($L$), while human satisfaction with personalization is assessed using the takeover rate ($R$).
Driving Performance The overall driving performance of Talk2Drive is reflected by a driving score ($S$), which is a weighted sum of four sub-scores: Time to Collision score ($S_{\tau}$), speed variance ($S_{\sigma}$), mean absolute acceleration ($S_{|\bar{A}|}$), and mean absolute jerk ($S_{|\bar{J}|}$).
$S=\sum{w_{k}\cdot S_{k}}$ (58)
For Time to Collision $\tau$, the critical threshold $\tau_{c}$ is based on human reaction time to take over and brake. A TTC greater than 1.5 seconds is often considered safe as the driver will have enough time to react to avoid rear-end collisions [van1994time]. Therefore, it can be considered as a hard threshold where any value greater than $\tau_{c}$ has a score of zero.
$S_{\tau}=\begin{cases}0,&\text{if }\tau_{min}<\tau_{c}\\ 100,&\text{if }\tau_{min}\geq\tau_{c}\end{cases}$ (59)
where $\tau_{c}$ is the critical threshold. We chose the value of 1.5 based on existing research results on Time to Collision value [van1994time]. For other metrics like speed variance, acceleration, and jerk, the thresholds are related to individual perception and would vary in different driving scenarios. Therefore, we define the score to quantify the corresponding performance, with a higher score indicating better performance relative to the baseline value. For example, the sub-score for mean absolute jerk $|\bar{J}|$ can be denoted as:
$S_{|\bar{J}|}=100-\gamma\cdot\frac{|\bar{J}|}{|\bar{J}|_{\text{baseline}}}$ (60)
@@ IV-C3 Evaluation Metrics
LLMs’ Latency for Time Efficiency Latency $L$ measures the response time of the LLM, which is important for time-sensitive vehicle applications. To measure latency, we calculate the time difference between the moment a command is sent to the LLMs and when the LLMs return the LMPs.
@@
Takeover Rate for Personalization The frequency of manual interventions indicates the model’s ability to adapt to personalized preferences from different humans [wang2022gaussian].
IV-D Experiment Results
In the experiments, all participants are required to issue verbal commands to the systems, which will be either baselines or Talk2Drive systems chosen randomly. To avoid bias, the participants will be unaware of which system is in use. They can decide when to take over the driving system, and the takeover rate will be recorded. Takeovers in this context are noted when humans find the current autonomous system unsatisfactory. Additionally, all driving data will be logged, and the ping latency for our experiments is at 200$\sim$400 ms.
IV-D1 The Validation of Driving Performance of Talk2Drive
@@
To validate the driving performance of the Talk2Drive system, we conduct comprehensive experiments to compare the driving performances between the baseline and the Talk2Drive through the metrics defined in Sec. IV-C3. As discussed earlier, the baseline values are from the average data of human drivers. The average driving performance for each scenario within each command directness is shown in Tab. VI. Note that for each expected driving behavior, the driving pattern is distinct. Therefore, a baseline is created for every driving behavior separately.
@@
The driving score is the weighted sum of all the safety and comfort metrics, reflecting the overall driving performance. The driving scores are calculated and recorded in Tab. VI, with green texts highlighting scores higher than the baseline of the same driving behavior and red texts showing the lower scores. As shown in Tab. VI, regardless of the command’s directness or the human’s behavior, nearly all driving scores are better than the baseline, showing that our framework consistently provides relatively safe and comfortable driving experiences for humans in all scenarios.
@@
The latency of the internet remains mostly consistent (around 300ms) under the same network conditions. The empirical results from our tests reveal that the latency for the most responsive LLM falls between approximately 1.2 to 1.8 seconds. This finding suggests considerable potential for integration into certain real-time applications where such response times are within acceptable limits. Specifically, LLMs demonstrate strong suitability for making high-level decisions or modifying driving behaviors in response to human commands. These scenarios require timely responses but do not typically demand extremely low latency.
IV-D2 The Improvement on Personalization Performance

### v4

@@ IV-D Summary
These findings, particularly the non-zero collision rates observed even in our best-performing models like the GPT-4 with human feedback (0.9% in LaMPilot-Bench), highlight a critical safety gap. While these models outperform rule-based methods in task completion, any persistent collision rate, no matter how small, is unacceptable for deployment in real-world, safety-critical autonomous driving.
This result strongly implies that the current generation of LLMs, when used in a zero-shot or few-shot manner, are not sufficiently reliable for direct, real-time motion control. The root cause, as discussed, often lies in the inherent latency and the non-real-time reasoning frequency of these large models. They struggle to account for the dynamic evolution of a driving scene between reasoning cycles, leading to failures in time-sensitive “corner cases.”
The safety implications of this are clear: LLM-based agents should not be granted direct, low-level control over a vehicle’s safety-critical functions. This limitation is a primary motivator for the architectural designs explored in our real-world experiments in Section V and VI. In those frameworks, the LLM is intentionally relegated to high-level, non-instantaneous tasks such as interpreting human preferences, personalizing driving style, or adjusting controller parameters while a separate, robust, low-latency control system maintains ultimate authority over real-time vehicle safety.
Additionally, NuPlanQA complements LaMPilot-Bench and CARLA Leaderboard by extending evaluation beyond closed-loop policy control to encompass perceptual understanding and reasoning in complex, real-world driving scenes. Together, these benchmarks form a holistic assessment pipeline for LLM4AD systems, spanning high-level perception and reasoning (NuPlanQA) to mid- and low-level decision execution (LaMPilot-Bench and CARLA).
V Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making
@@
V-A Talk2Drive Framework
@@
V-A1 Problem Statement
In Talk2Drive, we follow the concept we proposed in Sec. II. The model aims to translate verbal commands into executable control sequences for the vehicle. The cloud-based LLM acts as a translating function $f:I\rightarrow P$ that generates corresponding LMPs as the policy ($P$) for maneuvers.
For personalization, we need human feedback on execution $F(I,P)$ to evaluate if the generated policy addresses the preferences of the driver or the passengers. The extra memory module (see V-A5), which includes instructions, LMPs, and their evaluations, allows the LLM to learn user preferences. Therefore, there are two stages in the Talk2Drive workflow:
$\text{Execution}:$ $P\leftarrow f(I,S,C,H);$ (11)
$\text{Evaluation}:$ $H\leftarrow\left[I,P,F(I,P)\right].$
V-A2 Command Translation and Contextual Data Integration
$I\to\text{Could you drive more conservatively?}$ (12)
Simultaneously, LLMs access additional cloud-based real-time environment data, including weather updates, traffic conditions, and local traffic rules. In our product, LLMs can be informed by the weather information through Openweather API [openweather], the map information (such as road type and speed limits) through OpenStreetMap API [OpenStreetMap], and traffic information through TomTom API [tomtom]. We use a predefined structured language generator for both contextual data ($C$) and system messages ($S$), and our system will supply the appropriate values (the red contents) to the generator based on the context information, as shown below:
$C\to$ (13)
@@ V-A2 Command Translation and Contextual Data Integration
V-A3 Processing and Reasoning with LLMs
@@
{You are an autonomous vehicle with Adaptive cruisecontrol (ACC) and Lane Keeping Assist (LKA) alwaysenabled.You are using Pure Pursuit Controller to do the waypointfollowing.Here are some examples of how you need to react.Query: You drive too aggressively.Thought: The drivers think I drive too fast which looksaggressive and the drivers do not ask me to change lanes,so I need to slow down my speed.Action: $\left\{\begin{array}[]{l}\text${You are an autonomous vehicle with Adaptive cruise}\\ $\text${control (ACC) and Lane Keeping Assist (LKA) always}\\ $\text${enabled.}\\ $\text${You are using Pure Pursuit Controller to do the waypoint}\\ $\text${following.}\\ $\vdots\\$ $\text${Here are some examples of how you need to react.}\\ $\text${Query: You drive too aggressively.}\\ $\text${Thought: The drivers think I drive too fast which looks}\\ $\text${aggressive and the drivers do not ask me to change lanes,}\\ $\text${so I need to slow down my speed.}\\ $\text${Action: ...}\\ $\vdots\end{array}\right$. (34)
Additionally, the LLMs also access the memory module, a repository of historical interactions ($H$), to consider the human’s past behaviors and preferences. More details about the memory module are in Sec. V-A5.
V-A4 Actionable Code Generation
@@
[…]
The generated LMPs ($P$) are then sent back from the cloud to the vehicle’s electronic control unit (ECU), where they are executed. Additionally, we also set two kinds of safety checks for the generated LMPs. The first check involves verifying the format of the generated code. If the code does not conform to a valid format, our system will not execute or respond to the code. The second safety check is parameter verification, which assesses the feasibility and safety of the parameters within the context of the given policy, preventing the execution of potentially risky policies. For example, if the generated code instructs our agent to move to the left lane while the ego agent is already in the leftmost lane, our system will identify this abnormality and prevent the execution.
V-A5 Memory Module and Personalization
As described in Sec. II, the memory module is used to store the historical interactions ($H$) between humans and vehicles to provide a personalized driving experience. One important difference to note about this specific system, compared to the general framework, is that within this system, human feedbacks $F$ are archived within the historical memory, and the history memory $H$ includes three key linked components: the past commands ($I$) issued by the human, the corresponding LMP ($P$) generated by the LLMs in response, and the humans’ subjective evaluations or feedback ($F$) on the system’s performance for each command. Each interaction between the human and the vehicle is recorded and saved into a memory module in a text format within the ECU. This historical data in the memory module is updated after every trip. The example contents in the memory module are shown in Eq. 41:
@@ V-A5 Memory Module and Personalization
[…]
Figure 7: The flowchart of Talk2Drive. After the speech recognition module detects the keyword “command,” the inputs ($I,C,S,H$) are sent to the LLM. Then, the LLM generates corresponding LMPs to be executed by the ECU. If the speech recognition module detects the keyword “evaluate,” the system receives human feedback ($F$), and both $F$ and its corresponding II and $P$ are updated in the memory module.
@@ V-B Autonomous Driving System
V-C Experiment Setup
V-C1 Scenarios Setup and Participants
Our experimental trials include three different scenarios: highway, intersection, and parking lot.  [^2] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. The field experiments for the highway and intersection scenarios are conducted at a proving ground in Columbus, IN, USA, to validate the efficacy of the proposed Talk2Drive framework. The highway scenario involves a three-lane highway, while the intersection contains a two-way junction. Additionally, the parking lot scenario is evaluated at the North Stadium Parking Lot in West Lafayette, IN, USA. The overview visualization and statistics of the test tracks are shown in Fig. 9, while the visualization for the experiments is shown in Fig. 10. In our study, we recruit a diverse range of drivers, including seven individuals of different genders (61.4% male, 28.6% female), ages (mean=26.71, std=4.11), and driving experiences(mean=6.79, std=5.08). All participants have valid driving licenses.
V-C2 Input Instructions
@@
Given that conventionally indirect strategies primarily modify the politeness level of command of direct strategies, we instead divide the non-conventionally indirect requests into two categories based on the strength of the hints. We define our levels of directness in Tab. VII as our way to classify varying degrees of implicitness in the commands. To gather data, we request our test drivers to generate commands based on their normal speaking preferences and subsequently categorize their commands into our defined levels. Examples of the commands generated are also presented in Tab. VII.
TABLE VII: Level of Command Directness and Examples
@@
TABLE VIII: Driving Performance Validation
Driving
Secnario
Expected Driving
Behavior
Command
Directness
@@
V-C3 Evaluation Metrics
Our evaluation framework for autonomous vehicles includes driving performance, time efficiency, and personalization. We analyze Talk2Drive’s driving performance in terms of safety, using time-to-collision ($\tau$) and speed variance ($\sigma_{0}$), and comfort, using mean absolute acceleration ($|\bar{a}|$) and mean absolute jerk ($|\bar{J}|$). Time efficiency is measured by the LLM latency ($Lcy$), while human satisfaction with personalization is assessed using the takeover rate ($R$).
Driving Performance: The overall driving performance of Talk2Drive is reflected by a driving score ($Score$), which is a weighted sum of four sub-scores: TTC ($Score_{\tau}$), speed variance ($Score_{\sigma}$), mean absolute acceleration ($Score_{|\bar{a}|}$), and mean absolute jerk ($Score_{|\bar{J}|}$).
$Score=\sum{w_{k}\cdot Score_{k}}$ (60)
For TTC $\tau$, the critical threshold $\tau_{c}$ is based on human reaction time to take over and brake. A TTC greater than 1.5 seconds is often considered safe as the driver will have enough time to react to avoid rear-end collisions [van1994time]. Therefore, it can be considered as a hard threshold where any value greater than $\tau_{c}$ has a score of zero.
$Score_{\tau}=\begin{cases}0,&\text{if }\tau_{min}<\tau_{c}\\ 100,&\text{if }\tau_{min}\geq\tau_{c}\end{cases}$ (61)
where $\tau_{c}$ is the critical threshold. While 2.0 s is a common simulation margin, a 1.5 s threshold was selected for this real-world test based on existing research results on TTC value [van1994time]. For other metrics like speed variance, acceleration, and jerk, the thresholds are related to individual perception and would vary in different driving scenarios. Therefore, we define the score to quantify the corresponding performance, with a higher score indicating better performance relative to the baseline value. For example, the sub-score for mean absolute jerk $|\bar{J}|$ can be denoted as:
$Score_{|\bar{J}|}=100-\gamma\cdot\frac{|\bar{J}|}{|\bar{J}|_{\text{baseline}}}$ (62)
@@ V-C3 Evaluation Metrics
LLMs’ Latency for Time Efficiency: Latency $Lcy$ measures the response time of the LLM, which is important for time-sensitive vehicle applications. To measure latency, we calculate the time difference between the moment a command is sent to the LLMs and when the LLMs return the LMPs.
@@
Takeover Rate for Personalization: The frequency of manual interventions indicates the model’s ability to adapt to personalized preferences from different humans [wang2022gaussian].
V-D Experiment Results
[…]
V-D1 The Validation of Driving Performance of Talk2Drive
@@
To validate the driving performance of the Talk2Drive system, we conduct comprehensive experiments to compare the driving performances between the baseline and the Talk2Drive through the metrics defined in Sec. V-C3. As discussed earlier, the baseline values are from the average data of human drivers. The average driving performance for each scenario within each command directness is shown in Tab. VIII. Note that for each expected driving behavior, the driving pattern is distinct. Therefore, a baseline is created for every driving behavior separately.
@@
The driving score is the weighted sum of all the safety and comfort metrics, reflecting the overall driving performance. The driving scores are calculated and recorded in Tab. VIII, with green texts highlighting scores higher than the baseline of the same driving behavior and red texts showing the lower scores. As shown in Tab. VIII, regardless of the command’s directness or the human’s behavior, nearly all driving scores are better than the baseline, showing that our framework consistently provides relatively safe and comfortable driving experiences for humans in all scenarios.
@@
The latency of the internet remains mostly consistent (around 300 ms) under the same network conditions. The empirical results from our tests reveal that the latency for the most responsive LLM falls between approximately 1.2 to 1.8 seconds. While the observed 1.2 s to 1.8 s latency is acceptable for high-level, non-urgent decisions (e.g., responding to “drive more conservatively”), we must emphasize that this latency makes the cloud-based LLM entirely unsuitable for any safety-critical, time-sensitive maneuvers (e.g., emergency braking or collision avoidance.) This finding suggests considerable potential for integration into certain real-time applications where such response times are within acceptable limits. Specifically, LLMs demonstrate strong suitability for making high-level decisions or modifying driving behaviors in response to human commands. These scenarios require timely responses but do not typically demand extremely low latency.
Additionally, the Talk2Drive framework is designed to guarantee safety by architecturally separating the high-latency LLM from the low-latency, real-time vehicle control. As shown in Fig. 6, the LLM’s role is strictly limited to that of a high-level parameter adjuster. It does not engage in direct, real-time driving; instead, it generates LMPs that modify configuration parameters (e.g., target velocity, look-ahead distance) for the underlying, mature autonomous driving software such as Autoware, which handles the actual trajectory following and control. Furthermore, as detailed before, the system does not execute LLM-generated LMPs blindly. Every generated command is passed through a two-stage safety-check mechanism before execution. This includes: (1) a format check to ensure the code is valid and (2) a parameter verification step to assess the feasibility and safety of the proposed values. For example, any generated parameter that violates predefined safety bounds (e.g., a target velocity above the road’s speed limit) or is contextually impossible (e.g., an instruction to “move left” when the vehicle is already in the leftmost lane) is automatically identified and blocked from execution. This rule-based safety layer ensures that the LLM can only influence the vehicle’s driving style within a predefined, safe operational envelope, while all safety-critical responses remain the responsibility of the low-latency on-board controllers.
V-D2 The Improvement on Personalization Performance